The maximum amount of barium carbonate that will dissolve in a 0.247 M barium nitrate solution is ____M.

ksp BaCO3:8.1x-9
Answer #1


Ba(NO3)2 here is Strong electrolyte
It will dissociate completely to give [Ba2+] = 0.247 M
At equilibrium:
BaCO3     <---->     Ba2+     +          CO32-  
                     0.247 +s           s       
Ksp = [Ba2+][CO32-]
8.1*10^-9=(0.247 + s)*(s)
Since Ksp is small, s can be ignored as compared to 0.247
Above expression thus becomes:
8.1*10^-9=(0.247)*(s)
8.1*10^-9= 0.247 * 1(s)^1
s = 3.279*10^-8 M
Answer: 3.3*10^-8 M